Suppose a particle P is moving in the plane so that its coordinates are given by P(x, y),
where x = 4 cos 2t, y = 7 sin 2t.
(i) By finding a, b ∈ R such that x^2/a^2+y^2/b^2= 1, show that P is travelling on an elliptical path.
(ii) Let L(t) be the distance from P to the origin. Obtain an expression for L(t).
(iii) How fast is the distance between P and the origin changing when t = π/8?
1
Expert's answer
2020-06-29T18:23:15-0400

As "\sin^22t+\cos^22t=1" hence

(i) "\frac{x^2}{4^2}+\frac{y^2}{7^2}=1" is ellipse,where "a=4,b=7"

(ii) "L(t)=\sqrt{16\cos^22t+49\sin^22t}"

(iii) "L'(t)=\frac{66\cos2t\sin2t}{\sqrt{16\cos^22t+49\sin^22t}}" then "L'(\pi/8)\approx5.8" is the rate of change of distance between P and the origin